# Retinomorphic Machine Vision in a Network Laser

### v3

@@ Introduction
We now examine experimental feature detection. Figure 2g) shows six experimentally measured feature maps of the ‘CL’ image with colours corresponding to the modes shown in figure 1 and a composite feature map shown in figure 2h). Parallel feature detection of multiple image features is observed experimentally, with close correspondence to the netSALT simulated behaviour. Supplementary figures S6, S7 and S8 examine the feature detection performed by ten lasing modes across a range of test images. To compare our physically-detected feature maps against linear matrix kernel filters as used in software CNNs, we extract effective kernel filters for our modes via least-squares fitting between input images and the experimentally measured mode intensity (details in supplementary information). The extracted effective filters are shown in fig. 2i) with further details in supplementary fig. S9. Notably, both positive and negative kernel weights are observed - generating bipolar weights is highly challenging in neuromorphic hardware and an attractive benefit of the photonic lateral inhibition processes in our system which enable mode amplitude to both increase and decrease in response to specific image features. The $R^{2}$ values below each effective kernel indicate how well the linear fitting process captures our experimental results. When performing the same kernel fitting process on arbitrary linear software kernel filters, the process performs perfectly with $R^{2}$ values of 1.0. The kernels extracted from our experimental data have $R^{2}$ values from 0.628-0.946 (see fig. S9), suggesting that aspects of the physical feature detection process cannot be fully captured by linear kernel filters.
The experimental feature detection capability is proven for a higher-resolution (300$\times$200 pixel) photographic image of a street scene, shown in figures 2j,k). Separate lasing mode feature maps are shown for all test images in supplementary figures S10, S11 and S12.
Figure 3: Multilayer image classification. a) Feature detection: Input images are fractured into 4x4 pixel windows which are illuminated on the InP network in a sequential raster scan. The lasing amplitude of ten selected modes are used to perform spectrally-multiplexed feature detection and generate ten feature maps (three shown here, all ten shown in SI). b) Classification: Detected feature maps are then illuminated onto the network as whole images in a single illumination step with no window scanning. Output spectra are recorded for each feature map, giving ten spectra per image. c) The ten spectra are then combined in parallel, and a single logistic regression step is applied during training to generate a set of weights, one per hyperspectral channel (36813 channels after feature selection for MNIST-digits). For testing, no further regression is applied - each hyperspectral channel is multiplied by the fixed weights for each class obtained during training then summed. The class receiving the highest sum is then given as the prediction. A test accuracy of 98.05% is achieved for MNIST-digits. d) The same architecture is employed for the Fashion-MNIST task. Two input images are shown with six feature maps, colours corresponding to the modes shown in figures 1 and 2. e) A test accuracy of 87.85% is achieved for Fashion-MNIST. f,g) Training length vs. test classification accuracy comparison for MNIST digits (f) and Fashion-MNIST (g), with a zoomed-in plot of classification performance at shorter training sets for each task. The photonic convolutional network is compared against four software approaches: Logistic regression performed on the raw image data, a multi-layer perceptron, a LeNet-5 software CNN and a larger CNN (network details in text).
Multi-layer convolutional image classification
The feature maps detected by the network can be leveraged for machine vision tasks such as image classification, effectively providing an initial convolutional processing layer. Here we demonstrate the ability of the system to classify images, benchmarking classification via the MNIST digits and Fashion-MNIST tasks. For each task, we experimentally process 40,000 training images with a test set of 4000 images and training sets of 125 to 36,000 images.
For image classification we employ a multi-layer architecture, depicted schematically in figure 3a-c): an initial convolutional layer which performs feature detection as described above (fig. 3a), then a second layer where the detected feature maps are illuminated on the network as a single whole image (fig. 3b) with a logistic regression step applied to the spectral output of the second layer to generate classification results (fig. 3c).
@@
Ten spectral regions performing feature detection were selected, producing ten feature maps per input image. Three feature maps are shown schematically in figure 3b), with all feature maps for a range of MNIST digits and Fashion-MNIST shown in supplementary figures S13 and S14. These feature maps are then re-illuminated on the InP network as whole images, illustrated in the experimental camera image in figure 3b). The hyperspectral output from the second-layer feature map illuminations are combined in parallel and a single logistic regression step is applied in software to produce classification results. The same approach is employed for both MNIST digits (fig. 3a-c) and Fashion-MNIST (fig. 3d,e) tasks. We achieve an accuracy of 98.05% and 87.85% for MNIST digits and Fashion-MNIST respectively, further details of the regression process including feature selection are given in the Methods.
As a control, we perform the same final software logistic regression step directly on the input image data, bypassing the optical system. Substantially lower accuracies are observed of 92.10% and 82.29% for the MNIST digits and Fashion-MNIST tasks respectively.
@@
Beyond comparisons against logistic regression, figures 3f,g) show a comparison of classification accuracy vs. training dataset length for a range of software neural networks. Five approaches are compared: the photonic network, software logistic regression, a software multi-layer perceptron, and two software CNNs: one following the LeNet5 architecture, and a larger CNN with more filters and neurons. Full details of the comparison networks are provided in the methods. Training dataset length is varied from 125 to 36,000 images, with a fixed 4000 image test set. We also benchmarking against quasi-isomorphic software networks, comprising an initial convolutional filter layer of 10$\times$ 4$\times$4 filters followed by 30,000 randomly weighted ReLu neurons and a final logistic regression layer to mirror the structure of the experimental system. The results are shown in the supplementary information for two cases, one with random kernel filter weights and one using the kernel filter weights fitted from experimental data as shown in figure 2i).
At shorter training set lengths, the photonic network achieves higher accuracy than the software benchmarks. The photonic network scores highest for training sets below 5000 (digits) and 15,000 (fashion) images, demonstrating the learning strengths of the photonic system when working with short training datasets.
As one of the key proposed applications for neuromorphic computing hardware is ‘edge computing’ where remotely deployed hardware performs computation directly at the point of sensing, it is highly advantageous to perform well with short training sets. As edge-computing applications often demand that a system is able to adapt on-the-fly to changing environmental conditions and perform computation on remotely collected data without access to large offline training sets, the ability to perform well on short training sets is a strong benefit of this system.
@@ Conclusions
We have demonstrated a photonic machine vision platform based on a network laser. The system is capable of performing parallel feature detection, nonlinear activation and image classification in a single 150 $\upmu$m on-chip InP network with encouraging classification results, particularly on short datasets.
As the lasing modes responsible for feature detection spatially overlap, within the same InP network we are able to provide compact feature detection of multiple image features. In contrast, existing physical convolutional hardware schemes typically detect one feature per repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected. While we currently experimentally detect ten parallel image features, we find up to forty in simulation and the number of detected features is projected to increase as we deepen understanding of the links between network topology and feature sensitivity.
A key result of this work is demonstrating the importance of including both excitatory and inhibitory interactions in neuromorphic hardware. In biological systems, both of these interactions and their interplay are crucial to providing neuronal functionality[18, 16, 15], and existing physical neuromorphic hardware platforms often focus on excitatory signals. By engineering inhibitory nonlinear mode-competition between network lasing modes, we demonstrate the functional benefits of bio-inspired inhibitory dynamics.
As we develop finer control over the network dynamics, one can envision networks engineered to detect arbitrary specific features or provide reconfigurable detection via altering network dynamics via electrical gating or selective illumination of network sub-regions. More broadly, our approach shows the benefits of designing physical computing systems with intrinsic physical nonlinearities and a dynamic balance of competing excitatory and inhibitory interactions.
@@
RS, KM, HS, AF, DS and JD designed and developed the network laser concept.
@@
JCG, KS and TF performed machine learning analysis.
@@
AF designed and produced the figures with contributions from JD and WKN.
@@ Acknowledgements
We thank David Mack for excellent laboratory management and Steve Cussell and Dave Bowler for excellent technical workshop services at Imperial College London.
@@ Methods
The InP networks are fabricated on a Si chip with a 2 $\mu$m buffer layer of SiO2 (SOI chip) according to [30]. To obtain a defect-free InP layer on SOI, we used a Direct Wafer Bonding approach where the 127 nm thick InP layer is grown on a sacrificial III-V semiconductor wafer, with lattice-constant matched growth. This wafer is then annealed onto the SOI wafer and the sacrificial wafer removed by wet etching, resulting in a InP on SOI wafer that is cleaved into chips. The network structures are then patterned into a spin-coated hydrogen silsesquioxane (HSQ) mask with electron beam lithography. The pattern is transferred into the InP layer using Inductively Coupled Plasma (ICP) Reactive Ion Etching (RIE)[30] using a cyclic two-stage process: a CH4 and H2 plasma stage for etching, followed by O2 plasma stage for cleaning. To reduce surface state recombination losses and oxidation, the structures are passivated with phosphoric acid and coated with 3 nm of Al2O3 using Atomic Layer Deposition (ALD) [34].
@@
The network was designed by using a random point pattern on a plane and generating a Voronoi diagram[29]. The network is cropped to a circular area with 150 $\upmu$m diameter, and the density of the point patterns is varied to generate waveguide graph-edges with 350 nm width and an average length of 5 $\upmu$m.
@@
The lasing modes and intensities of the network are modelled using netSALT[25]. The designed network is represented as a graph, with the edges subdivided to overlap with an 8x8 grid to correspond to the kernel. When the network is unpumped, each edge has a complex refractive index n = 3.4 + 0.001i. The edges are further assumed to have a Lorentzian gain profile centered at normalised frequency $k_{a}=7$ with a width $\gamma_{a}=0.35$, and are pumped up to a normalised power $D_{0}=0.5$ (For a detailed physical explanation of the parameters, see Ref. 25). A separate simulation is performed for each 8x8 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane $k$, and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im$(k)=0$], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.
@@
We benchmark our system against four software approaches for the MNIST and Fashion-MNIST tasks: logisitic regression, a multi-layer perceptron and a convolutional neural network (LeNet5) and a larger convolutional neural network model with many more filters and trainable parameters.

### v4

@@ Introduction
We now examine experimental feature detection. Figure 2g) shows six experimentally measured feature maps of the ‘CL’ image with colours corresponding to the modes shown in figure 1 and a composite feature map shown in figure 2h). Parallel feature detection of multiple image features is observed experimentally, with close correspondence to the netSALT simulated behaviour. Supplementary figures examine the feature detection performed by ten lasing modes across a range of test images. To compare our physically-detected feature maps against linear matrix kernel filters as used in software CNNs, we extract effective kernel filters for our modes via least-squares fitting between input images and the experimentally measured mode intensity (details in supplementary information). The extracted effective filters are shown in fig. 2i) with further details in supplementary information. Notably, both positive and negative kernel weights are observed - generating bipolar weights is highly challenging in neuromorphic hardware. This is an attractive benefit of the photonic lateral inhibition processes in our system, which enable mode amplitudes to both increase and decrease in response to specific image features. The $R^{2}$ values below each effective kernel indicate how well the linear fitting process captures our experimental results. When performing the same kernel fitting process on arbitrary linear software kernel filters, the process performs perfectly with $R^{2}$ values of 1.0. The kernels extracted from our experimental data have $R^{2}$ values from 0.628-0.946 (see SI), suggesting that aspects of the physical feature detection process cannot be fully captured by linear kernel filters.
We extend experimental demonstration of feature detection to high-resolution colour images, showing a flexible range of kernel sizes using the popular ‘mandrill’ test image (512$\times$512 pixels, not a baboon as often thought!)[39]. Figure 2j) shows the input image, which is separated into RGB channels and illuminated onto the network at a range of kernel sizes to demonstrate the capability of our system to detect features with arbitrary kernel window size. This flexibility is typically not available in the more common matrix-multiplication style feature detection schemes where kernel size is hard-coded at fabrication stage[34]. Figure 2k) shows experimentally detected feature maps for 4$\times$4, 7$\times$7 and 11$\times$11 kernels, with the R/G/B coloured features detected from the corresponding colour channel (the response of all ten feature-detecting modes are summed to produce a single feature map for each colour channel here). The ability to tune feature detection sensitivity to higher or lower spatial frequency is shown in our feature maps produced by different kernel sizes, the 4$\times$4 kernel picks up high-frequency details in the mandrill fur around image corners, whereas the 11$\times$11 kernel exhibits weaker response to these. Increasingly, powerful modern software CNNs such as EfficientNet[40, 41] employ a range of filter sizes, demonstrating the utility and attraction of our ability to freely select filter size. Separate lasing mode feature maps are shown for a wider range of test images in supplementary information.
Network Laser Image Classification
Here we demonstrate the ability of the network laser to classify images. The network can be operated in a single-layer scheme, directly projecting images onto the network (fig. 3a), or a two-layer scheme with an initial feature-detection layer (fig. 3b), followed by projecting the feature-maps onto the network (fig. 3c). In both cases, a logistic regression step follows the final layer of hyperspectral output to generate classification predictions.
We perform classification on three tasks: MNIST digits, Fashion-MNIST, and the challenging low-data BreaKHis breast cancer diagnosis[42] (colour histopathology images, 400$\times$ magnification, 700$\times$460 resolution). Figures 3d,e,f) show example images for each task, with coloured lines illustrating feature maps detected by the network. For MNIST digits and Fashion, we experimentally process 40,000 training images with a test set of 4000 images and training sets of 125 to 36,000 images. For BreaKHis we process 1693 images, with a test set of 250 and training sets of 10 to 1,443 images. BreaKHis contains a roughly 2:1 ratio of malignant to benign images, with this class-imbalance increasing task difficulty.
Single-layer classification (fig. 3a) works by projecting whole images onto the network, measuring the hyperspectrum, then applying logistic regression on the training set hyperspectra to learn a fixed weight (per image class) for each channel of the hyperspectral output. Once the system is trained, unseen images are classified by performing a weighted sum over the hyperspectral channel amplitudes and learned weights for each class. The class with the largest weighted sum is returned as the network prediction. The single-layer scheme obtains accuracy of 96.03% on MNIST digits, 85.18% on Fashion MNIST, and 86.59% on BreaKHis cancer diagnosis, with measurement throughput up to 100 Hz.
@@
Multi-layer classification begins with feature detection (fig. 3b). Ten spectral regions performing feature detection were selected, producing ten feature maps per input image. Regions returning clear edge-detection were hand selected, with future improvement available by training the feature selection process. Three feature maps are shown schematically in figure 3b), with all feature maps for a range of MNIST digits and Fashion-MNIST shown in supplementary information. These feature maps are then re-illuminated on the InP network as whole images and their corresponding hyperspectra recorded, illustrated in figure 3c). The hyperspectral output from the second-layer feature map illuminations are combined in parallel and a single logistic regression step is applied to produce classification results. The multi-layer scheme obtains accuracy of 98.05% on MNIST digits, 87.85% on Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis. Further details of the regression process are given in the Methods.
Figure 4: Biomedical Image Segmentation and Class-Imbalanced Diagnostic Classification on the HAM10,000/ISIC 2018 dataset. a) Example images of the seven skin lesion classes contained in the dataset. Images are complex with different instances of the same class looking very different, hard to diagnose without expert medical training.
b) Pie-chart illustrating the heavy class-imablance in the dataset.
c) Class-balanced accuracy vs. training set length, comparing photonic network against software benchmarks. Test set is fixed at 1015 images. Class-balanced accuracy penalises incorrect classification on rarer classes more heavily, ensuring performance is not dominated by majority classes. Photonic network scores 53.7% max accuracy, beyond software models.
d) Segmentation scores: DICE, Jaccard, and pixel accuracy of photonic network and software benchmarks. e-g) Image segmentation results, comparing photonic network performance against software benchmarks for three example test set images. Ground-truth masks (yellow line) generated by a human medical expert[43] are used for supervised training, and assessment of network performance on test images. Networks generate predicted segmentation masks (blue line) around where they think the lesion boundaries are. Green regions indicate correct segmentation where the network prediction agrees with the ground truth, red regions show combined false positive & negative errors.
@@
Network laser performance relative to software benchmarks is assessed in figures 3g,h,i). Classification accuracy vs. training set length is assessed for single- and multi-layer network laser schemes against a range of software neural networks. Five software approaches are considered: logistic regression on the raw images (a direct control, effectively bypassing our physical system), a multi-layer perceptron (300$\times$100 neurons), a LeNet5 CNN (22 filters), a larger CNN (96 filters), and the much more powerful ‘EfficientNetV2-B0’[40, 41] CNN (3040 filters, 7.9 million parameters) using a transfer learning approach - initialised with pre-trained ImageNet weights, which are then updated during training. Full details of the software networks are provided in the Methods. We also benchmark against a ‘quasi-isomorphic’ software network: an initial convolutional filter layer of 10$\times$ 4$\times$4 filters followed by 30,000 randomly weighted ReLu neurons and a final logistic regression layer, mirroring the structure of our system. Results are shown in supplementary information.
Few-shot/low-data performance The photonic network laser excels when training data is scarce, a regime where conventional AI can struggle. The network outperforms all considered software models on training sets smaller than 5,000 images (digits), 15,000 images (fashion), and across all training lengths on the harder biomedical BreaKHis task — including the powerful EfficientNet CNN. We score 77.7% accuracy on BreaKHis from just 10 training images (5 examples per class), far higher than the software models. This highlights the effective learning provided by the nonlinear photonic dynamics and mode competition of the network laser. Here we compare performance when using only the raw training set. Software approaches can achieve higher scores on the BreaKHis task, but this requires costly extra steps including artificial augmentation of the training set, and extensive fine-tuning of model weights.
The origin of this strong few-shot performance may lie in the system’s high-dimensional, heterogeneous nonlinear responses. Recent theoretical work has shown that if a learning network is sufficiently ‘high dimensional’ (e.g. it has a large number of output features/channels with unique nonlinear responses), then the network may learn to generalise/accurately classify from very few training examples[11]. This behaviour is exhibited in our network: a broad range of unique nonlinear responses is provided by the many lasing modes, shown experimentally in figure 1g). The diverse nonlinearity in the ‘photonic neurons’ (our lasing modes) provide both high dimensionality and neuronal heterogeneity, which has been shown to promote robust learning and stronger task performance in models of biological neurons[22].
Combined Biomedical Image Segmentation and Class-Imbalanced Diagnostic Classification.
The neuromorphic capabilities of the photonic network are not restricted to feature detection and classification. Here, we demonstrate the ability of the network to spatially map regions of interest in biomedical images - drawing masks over skin lesions in colour photographic images from the HAM10000/ISIC 2018 dataset[44, 43] (fig. 4a, 10015 images of seven different classes of skin lesion). This task, termed segmentation, demonstrates that while the hyperspectral output of our network only contains information on the wavelength and one-dimensional spatial position of the lasing response, it is capable of physically encoding and processing complex two-dimensional spatial information - required for segmentation and other spatial processing. The task here is challenging: lesions are non-uniform in shape, a variety of colours and intensities, and many images contain darker non-lesion regions such as hairs, moles, and areas of uneven lighting. We combine two distinct tasks on the HAM10k dataset: class-imbalanced diagnostic classification, separating lesions into seven classes, and segmentation, drawing spatial masks to indicate lesion location. This two-fold processing has potential future impact in automated/semi-automated combined diagnosis and surgery systems. We need only input the HAM10k images to our photonic network once to achieve both tasks, simply retraining our computationally-cheap linear output regression layer to change between diagnosis and segmentation.
Class-imbalanced diagnosis Here we diagnose lesion images into seven classes. This is made highly challenging by the heavy class-imbalance of the dataset - illustrated in figure 4b). Machine learning typically struggles with class imbalance, tending to incorrectly classify rare classes as examples of the more common classes. Overcoming this requires costly techniques such as data augmentation, e.g. generating many distorted copies of training examples and using them to artificially extend the training set. The HAM10000 task is heavily imbalanced, with the ‘NV’ class representing 66.9% of the dataset, and four of the seven classes each representing 5.1% or less. Figure 4c) shows class-balanced classification accuracy against training set length (class-balanced accuracy penalises incorrect guesses on rare classes more heavily), with the photonic network scoring 53.7% max accuracy, outperforming all software benchmarks at all training lengths. This shows our system demonstrates strong advantages on both few-shot learning and heavily-imbalanced datasets - a hard problem.
Biomedical segmentation For segmentation, we use the same hyperspectral responses from the lesion images which we use for classification. Here, rather than using the diagnostic class as training/testing labels, and logistic regression as an output layer - we use the ground-truth segmentation maps as training/testing labels and ridge regression as an output layer (as the segmentation task has continuous output, rather than discrete labels). Otherwise the procedure follows that used for classification. Figure 4d) compares three segmentation metric scores across the test set: DICE (F1 score for segmentation), Jaccard (intersection over union) and pixel accuracy (correct pixels over incorrect). DICE and Jaccard provide best indication of segmentation quality, pixel accuracy scores tend to be distorted by whether the correct ground-truth mask occupies a small or large fraction of the total image. The photonic network scores highest on all metrics, highlighting the ability of the network to implement spatial processing.
Figures 4 e-g) show segmentation results for the photonic network and software benchmarks for three test images. In the top row (e), the photonic network accurately predicts the lesion location, ignoring the non-lesion mole, unlike the CNN and ridge regression software models. In the middle row (f) the photonic network outperforms software on a complex lesion shape with regions of varying intensity, and in the bottom row (g) the photonic network avoids incorrectly classifying the dark corner regions or hairs.
@@ Conclusions
This work introduces a new class of photonic hardware for compact, highly nonlinear neuromorphic processing which excels at few-shot learning — with potential impact on biomedical sensing, edge computing, and beyond. The system performs parallel feature detection, heterogenous nonlinear activation, and strong image classification and segmentation in a single 150 $\upmu$m on-chip InP network, outperforming powerful modern software CNNs such as EfficientNet particularly on short and imbalanced datasets.
We provide routes to solutions for both the physical nonlinearity and large footprint issues facing many existing photonic computing schemes. As the lasing modes responsible for feature detection spatially overlap, we provide compact feature detection of multiple image features in a micron-scale footprint. In contrast, existing physical convolutional hardware schemes typically detect one feature per physical repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected.
Strong learning in the few-shot/low-data regime with scarce training examples is a hard and open challenge in machine learning. The performance of our photonic network in this regime, including on hard biomedical tasks such as BreaKHis and HAM10k, provides valuable evidence that physics-based neuromorphic systems can handle complex, data-constrained learning tasks. These results bring physical neuromorphic schemes closer to offering viable alternatives to conventional software-based approaches. Our good performance here extends to heavily class-imbalanced datasets, which are similarly challenging for conventional software AI. One of the key proposed applications for neuromorphic computing hardware is ‘edge computing’ where remotely deployed hardware performs computation directly at the point of sensing, adapting on-the-fly to incoming data and changing environmental conditions without access to large offline/cloud-based training data. Here, it is highly advantageous to perform well with short and imbalanced training sets, as locally collected data will typically be far from the ideal large and class-balanced training sets where software AI is most comfortable.
A key result of this work is demonstrating the importance of including both excitatory and inhibitory interactions in neuromorphic hardware. In biological systems, both of these interactions and their interplay are crucial to providing neuronal functionality[19, 21, 17], and existing physical neuromorphic hardware platforms often focus solely on excitatory signals. By engineering inhibitory nonlinear mode-competition between network lasing modes, we demonstrate the functional benefits of bio-inspired inhibitory dynamics.
As we develop finer control over the network dynamics, one can envision networks engineered to detect arbitrary specific features or provide reconfigurable detection via altering network dynamics through electrical control or selective illumination of network sub-regions[26] to provide effective tunable physical weights in the input space[45]. More broadly, our approach shows the benefits of designing physical computing systems with diverse physical nonlinearities, and a balance of competing excitatory and inhibitory interactions. While our system currently exhibits strongest performance on harder, lower training data tasks, we foresee that future iterations will extend this performance to larger datasets via training of the physical dynamics and optimisation of the network graph topology.
@@
JCG, JD and TVR developed the feature detection scheme.
RS, KM, HS, J.D., AF, DS, and W.K.N. designed and developed the network laser concept.
@@
JCG, JD, KS and TF performed machine learning analysis.
@@
AF, WKN and JD designed and produced the figures.
@@ Acknowledgements
We thank David Mack for excellent laboratory management and Stephen Cussell and David Bowler for excellent technical workshop services at Imperial College London.
@@
Test image of ‘Mandrill’ was taken from the USC-SIPI image database, available at: https://sipi.usc.edu/database/database.php
@@ Methods
The InP networks are fabricated on a Si chip with a 2 $\mu$m buffer layer of SiO2 according to [33]. To obtain a defect-free InP layer on SOI, we used a Direct Wafer Bonding approach where the 127 nm thick InP layer is grown on a sacrificial III-V semiconductor wafer, with lattice-constant matched growth. This wafer is then annealed onto the SOI wafer and the sacrificial wafer removed by wet etching, resulting in a InP on SOI wafer that is cleaved into chips. The network structures are then patterned into a spin-coated hydrogen silsesquioxane (HSQ) mask with electron beam lithography. The pattern is transferred into the InP layer using Inductively Coupled Plasma (ICP) Reactive Ion Etching (RIE)[33] using a cyclic two-stage process: a CH4 and H2 plasma stage for etching, followed by O2 plasma stage for cleaning. To reduce surface state recombination losses and oxidation, the structures are passivated with phosphoric acid and coated with 3 nm of Al2O3 using Atomic Layer Deposition (ALD) [48].
@@
The network was designed by using a randomly-distributed point grid pattern on a plane and generating a Voronoi diagram[32]. The network is cropped to a circular area with 150 $\upmu$m diameter, and the density of the point patterns is varied to generate waveguide graph-edges with 350 nm width and an average length of 5 $\upmu$m.
@@
Throughput of images was run up to 100 Hz, with the limiting factor currently the camera detection. We have run experimental tests on a similar system up to 500 Hz. With the correct camera, throughputs of 1 kHz are readily feasible without modification to the scheme. Beyond this, the next limiting factor is the rate at which the DMD can display new images. Our current DMD has 10 kHz speed, and faster DMDs are available up into the MHz, with ongoing improvements in DMD technology likely to improve this in the near future. The repetition rate of our pulsed laser goes to 100 kHz, with MHz pulsed lasers or on-chip light sources available. The eventual fundamental limit is the 10-100 ps timescale of the photonic dynamics. Beyond this, more fundamental changes to the physical system/excitation scheme will be required.
A second 80 picosecond pulse pump laser was used to confirm that we are not limited to operating with femtosecond pulses, and can operate with a range of pump wavelengths. We employed the Picophotonics CP32 532 nm MOPA pump laser, which successfully pumps our InP network to lase. Experimental data confirming this (hyperspectra, LL plots) are shown in the supplementary information.
@@
The lasing modes and intensities of the network are modelled using netSALT[26]. The designed network is represented as a graph, with the edges subdivided to overlap with an 8$\times$8 grid to correspond to the kernel. When the network is unpumped, each edge has a complex refractive index n = 3.4 + 0.001i. The edges are further assumed to have a Lorentzian gain profile centered at normalised frequency $k_{a}=7$ with a width $\gamma_{a}=0.35$, and are pumped up to a normalised power $D_{0}=0.5$ (For a detailed physical explanation of the parameters, see Ref. 26). A separate simulation is performed for each 8$\times$8 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane $k$, and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im$(k)=0$], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.
@@
Regression process for the network laser image segmentation task
The process is nominally identical to the image classification workflow, save for these details on replacing logistic regression with ridge regression, and changing the ground truth from classification labels to segmentation maps:
Ridge regression is used instead of logistic regression (using sklearn’s built in ridge regression function) as we want continuous outputs here. Logistic regression returns discrete outputs, such as class-based classification predictions. Segmentation is performed by generating 2D maps of confidence/probability that a pixel is part of a lesion, hence ridge regression is used as the output layer. The confidence maps are then thresholded at 50% confidence to produce binary maps, and regions with higher than 50% are predicted to contain lesions. We experimented with varying thresholds and having this as an optimised hyperparameter, but the best performing threshold values were so close to 50% (plus or minus 1% typically) that we leave the threshold fixed at 50% to simplify the process.
The ground truth labels used for training and testing in the segmentation tasks are 2D maps produced by a human medical expert[43]. We train fixed weights on each hyperspectral output channel which generate segmentation maps based on network hyperspectra, trained by maximising DICE accuracy against the ground truth maps. We train on 9000 images, and test on 1015.
@@
We benchmark our system against five software approaches for the classification tasks: logistic regression, a multi-layer perceptron, a small convolutional neural network (LeNet-5), and a larger convolutional neural network model with more filters (96) and trainable parameters, and a far larger EfficientNetV2-B0 CNN (7.9 million parameters, 3040 filters) pre-trained using ImageNet weights and operated in a transfer learning scheme.